OpenStudy (anonymous):

How do I rewrite sin^(4)xtan^(2)x in terms of the first power of cosine?

OpenStudy (anonymous):

your hot and pretty

OpenStudy (anonymous):

answer me please

OpenStudy (anonymous):

\[\sin^4 x \tan^2x = \sin^4 x \frac {\sin^2x}{\cos^2x} = \frac {(\sin^2x)^3}{\cos^2x} =\frac {(1-\cos^2x)^3}{\cos^2x}\] Simplify it then the final answer is: \[-\cos^6x+3\cos^4x-3\cos^2x+1\]